OpenStudy (anonymous):

When two consecutive odd integers are multiplied together the answer is 4623. (a) If the smaller of the two integers is n, what is the larger? (b) Form a quadratic equation in n and solve it.

OpenStudy (anonymous):

b) so the quadratic is n(n+2)=4623. a) if the smaller is n then the larger is 4623/n

OpenStudy (anonymous):

The answers are 67 and 69 by the way.

OpenStudy (vishal_kothari):

it's right...

OpenStudy (anonymous):

yeah. correct. the same computation!!

OpenStudy (anonymous):

n and n+2 is the larger n(n+2)=4623 n^2 + 2n =4623 n^2 +2n - 4623 =0 n=-69 or n=67 n+2=-67 or n+2 = 69

OpenStudy (mathmate):

If it's just the answer you want, then 1. take the square-root, round to the nearest integer. 67.992 => 68 2. Add/subtract 1. (67, 69)